### v2

# FeTa: A ferrimagnetic quantum anomalous Hall insulator

Abstract
We theoretically propose that the van der Waals layered ternary transition metal chalcogenide FeTa ( S, Se, Te) is a new family of ferrimagnetic quantum anomalous Hall insulators with sizable bulk gap and Chern number . The magnetic ordering originates predominantly from the Fe atoms, where strong ferromagnetic exchange interactions between them induce magnetic moments on the Ta sites, yielding a collinear ferrimagnetic ground state. The large topological gap arises from the deep –-type band inversion between spin-down Ta and orbitals near the Fermi level—a mechanism unique to -orbital systems. Remarkably, the Curie temperature of monolayer FeTa is predicted to significantly exceed that of monolayer MnBi2Te4. Furthermore, both the Curie temperature and topological gap scale positively with the spin-orbit coupling strength of the electrons, suggesting a common physical origin. These findings, if realized experimentally, could open new avenues for the research and application of topological quantum physics.
Introduction—The discovery of the quantum anomalous Hall (QAH) effect marked a milestone in the study of topological quantum states [1, 2, 3, 4]. These states exhibit an insulating bulk characterized by a finite Chern number and support gapless chiral edge modes [5, 6, 7]. The QAH phase provides a versatile platform for realizing exotic phenomena such as Majorana edge modes [8, 9, 10, 11, 12] and the quantized topological magnetoelectric effect [13, 14, 15, 16, 17], with potential for dissipationless electronic applications [18, 19]. Experimental realizations of the QAH effect span a range of material systems [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34], yet all occur at liquid helium temperatures. This low critical temperature remains a major barrier to practical use, including in metrological standards [35, 36, 37]. Identifying new QAH insulators [38, 39, 40, 41, 42, 43, 44, 45, 46] with higher transition temperatures and larger bulk gaps is therefore a central challenge in topological materials research.
The QAH effect originates from spin-polarized band inversion [47], where achieving high transition temperatures and large bulk gaps necessitates a delicate interplay between two seemingly conflicting requirements: strong spin-orbit coupling (SOC) and stable ferromagnetism. From a material design perspective, strong SOC favors heavy elements, whereas ferromagnetism typically emerges in metallic systems composed of lighter element. This incompatibility has not been fully resolved in existing QAH systems. In magnetic topological insulators (TIs), the inhomogeneities drastically suppress the exchange gap by several orders of magnitude [48, 49, 50, 51, 52, 53]; moiré systems are intrinsically limited by their low energy scales [29, 30, 31]; and rhombohedral graphene suffers from a weak proximitized Ising SOC [54]. Therefore, finding stoichiometric 2D magnetic materials for the QAH effect—ideally in monolayer form with versatile tunability—is highly desired.
Materials | (Å) | (K) | (meV) | (meV) |
FeTaS2 | 3.74 | 725 | 363.3 | 11.2 |
FeTaSe2 | 3.80 | 770 | 401.7 | 13.0 |
FeTaTe2 | 3.91 | 560 | 218.7 | 18.5 |
A promising route toward high-temperature QAH insulators lies in leveraging correlated electrons, which can simultaneously support robust magnetism and nontrivial band topology. FeSe and its derivatives are particularly compelling in this context, as electron doping via lithium intercalation has revealed the coexistence of superconductivity, itinerant ferromagnetism, and topology [59, 60, 61, 62, 63, 64, 65]. Motivated by these developments, we perform density functional theory (DFT) calculations and tight-binding modeling to demonstrate that monolayer FeTa ( = S, Se, Te) emerges as a promising candidate for high-temperature QAH insulators. All compounds exhibit a ferrimagnetic ground state with Chern number and exceptionally large bulk gaps. Notably, the magnetic ordering originates primarily from the Fe atoms, while the topology arises from the spin-polarized band inversion involving the Ta and orbitals. This mechanism, where both magnetism and topology emerge cooperatively from -orbital physics, suggests a broadly applicable design principle for engineering QAH phases in correlated materials.
Structure and magnetic properties—The monolayer FeTa crystallizes in a tetragonal lattice structure with the space group - (No. 115). As shown in Fig. 1(a), each primitive cell includes three atomic layers—one FeTa and two layers—where each Fe or Ta atom is coordinated by four atoms forming a distorted edge-sharing tetrahedron. This structure can be viewed as a half-substituted derivative of Fe with space group , in which one of the two inequivalent Fe atoms in the unit cell is replaced by a Ta atom. Consequently, inversion symmetry () is broken, but the combined symmetry remains preserved [Fig. 1(b)]. The fully optimized lattice constants are listed in Table 1. The dynamical and thermal stability of monolayer FeTa are confirmed by first-principles phonon calculations and molecular dynamics simulations [73], respectively. The van der Waals nature of these materials suggests that monolayers could be experimentally exfoliated from bulk samples once synthesized.
The magnetic properties of monolayer FeTa are markedly altered by Ta substitution. First-principles calculations reveal that all compounds listed in Table 1 exhibit a robust collinear ferrimagnetic ground state with an out-of-plane easy axis, in contrast to monolayer FeTe, which is observed to have a stripe antiferromagnetic (AFM) ground state [74]. The total magnetic moment per unit cell is , where each Fe atom possesses sizable local moments (2.7), and Ta atom exhibits antiparallel moments (1.4). The magnetic ordering is predominantly governed by the Fe atoms, while the moments of Ta atoms are induced via strong molecular fields generated by neighboring Fe atoms [73]. This ferrimagnet effectively behaves as a ferromagnet, with the net magnetization aligned along the Fe moment direction. As a result, the system avoids a compensation temperature, consistent with Néel’s molecular field theory [75, 76].
To elucidate the origin of ferrimagnetism, we first analyze the orbital occupation of Fe atoms. In the tetrahedral crystal field, the Fe orbitals split into a lower-energy doublet and a higher-energy triplet [Fig. 1(d)]. The Fe atoms adopt a high-spin configuration: , yielding a spin magnetic moment of in accordance with Hund’s rule, consistent with our DFT results. The spin-up channel is fully occupied, while the partially filled spin-down channel is at the Fermi level. Consequently, direct electron hopping between neighboring Fe sites is allowed only when their magnetic moments are aligned parallel, as enforced by the Pauli exclusion principle. In addition, the superexchange interaction mediated by the ligands (with the nearly Fe––Fe bond) further stabilize ferromagnetic (FM) coupling between the Fe atoms, according to the Goodenough–Kanamori–Anderson rules [77]. This particular configuration of Fe orbitals in a tetrahedral crystal field leads to ultrastrong FM coupling between Fe atoms via both the FM kinetic exchange and superexchange mechanisms [40, 41, 45].
For the Ta atoms in FeTa, the crystal fields generated by the surrounding Fe atoms and ligand atoms compete, resulting in a relatively small splitting between the lower-energy doublet and higher-energy triplet of Ta orbitals. As a result, each Ta atom adopts an configuration, yielding a local magnetic moment of approximately . The reduced crystal field splitting enhances the FM - kinetic exchange, which dominates over the AFM - exchange between the neighboring Ta atoms [43, 44]. Namely, the coupling between Ta atoms is FM. Moreover, direct electron hopping between Fe atoms in the configuration and adjacent Ta atoms is energetically favorable when the latter adopt a configuration [Fig. 1(d)]. This facilitates an effective AFM exchange coupling between the Fe and Ta sublattices.
Collectively, these exchange interactions stabilize a ferrimagnetic ground state in monolayer FeTa, featuring strong FM coupling between Fe atoms and antiparallel alignment of the induced moments on Ta sites. The system exhibits exceptionally large magnetocrystalline anisotropy energy (MAE), exceeding that of known 2D magnets such as Fe3GeTe2 [78]. As listed in Table 1, the large positive MAE indicates a robust out-of-plane easy axis. Notably, the predicted Curie temperature for monolayer FeTa significantly surpasses that of MnBi2Te4.
Electronic structures— Figures 2(d) and 3(a) show the electronic structure of monolayer FeTaSe2 with and without SOC, respectively. In the absence of SOC [Fig. 3(a)], four spin-polarized Dirac points emerge along the -X(-Y) direction, which are protected by mirror symmetry (). The -orbitals-resolved analysis reveals that these Dirac cones originate from the band crossings between the Ta and orbitals. Upon inclusion of SOC, the out-of-plane ferrimagnetic ordering breaks both and , opening a finite gap at each Dirac point. Each gapped Dirac cone contributes a quantized Berry phase of , yielding a total Berry phase of due to the fourfold valley degeneracy. This results in a QAH conductance of [Fig. 2(f)], corresponding to a Chern number . The nontrivial topology is further confirmed by the presence of two chiral edge modes traversing the bulk gap in the edge local density of states [Fig. 2(e)]. The sign of aligns with the direction of the Fe minority spins. Substituting Se with isovalent chalcogen atoms S or Te generates similar band structure and preserves the overall band topology and Chern number, as illustrated in Figures 2(a-c) and 2(g-i). Notably, the Dirac gap increases from S to Te, reflecting the enhancement of SOC strength with heavier chalcogen elements.
The topological gap in FeTa exhibits a strong dependence on the spin orientation. As the magnetization direction rotates from to , and subsequently to , the bulk gap closes and reopens, signaling two topological phase transitions—from to , and then to [Fig. 3(b)]. Quantitatively, the gap follows an approximate relation , indicating that the topological gap is predominantly governed by the out-of-plane spin component of the Ta atoms via SOC. We point out that the large topological gap of the ground state still holds in the presence of magnon excitation, which is ensured by the large MAE [73]. The significant thermal spin fluctuation will decrease and the topological gap dramatically only when temperature is close to .
In FeTa, both the topology and magnetism originate from the orbitals of transition metal elements. To highlight their shared origin, Fig. 3(c) presents the dependence of the Dirac gap and Curie temperature on the SOC strength in monolayer FeTaSe2. As the SOC increases, both the topological gap and exhibit a clear positive correlation, indicating that SOC not only drives the nontrivial band topology but also enhances the magnetic stability. This concurrent enhancement emphasizes a key advantage of -orbital systems: the intrinsic coupling between magnetism and topology. Such synergy offers an effective strategy for realizing high-temperature QAH insulators with sizable topological gaps.
Model and origin of topology—While the band structure near the Fermi level might suggest that the band topology arises from an –-type band inversion at the X point—supported by the opposite eigenvalues of the irreducible representations X1 and X4 [Fig. 3(d)] and their potential to yield a high Chern number () via valley multiplicity—a more detailed orbital analysis reveals otherwise. The -orbital band projections show that X1 originates from the unoccupied Fe band, which does not participate in the band inversion. Instead, the nontrivial topology stems from the band inversion between the Ta and orbitals at the point, involving an exchange of irreducible representations and [73]. Given the considerable depth of this inversion, its associated features are not readily apparent near the Fermi level.
To elucidate the origin of the band topology, we construct a minimal tight-binding model that incorporates the Ta , and Fe orbitals. This model, detailed in the Supplemental Material [73], includes the nearest- and next-nearest-neighbor hopping terms as well as SOC. As shown in Figs. 4(a) and 4(b), it accurately reproduces the DFT band structure [Fig. 3(a)] and the irreducible representations at high-symmetry points (listed in Table 2). Importantly, while the model captures the apparent –-like features around the X point via inter-orbital hybridization, it demonstrates that the topological band structure is fundamentally driven by the –-type inversion between the Ta and orbitals at point.
We now analyze the origin of the large topological gap. The – band inversion at the point can be described by a minimal Hamiltonian,
where and are the effective mass of and bands near the point, respectively. characterizes the band inversion strength, typically several eV in FeTa. The parameters and are real: arises from inter-orbital overlap, while originates from SOC. Due to the angular momentum difference between the two orbitals, the inter-orbital couplings exhibit a quadratic -dependence. This model leads to a topological gap , which opens at finite away from the point due to orbital hybridization. In the case of FeTa with deep band inversion, where the relevant wave vectors lie near the X point, the gap is more accurately described by . This form highlights an enhancement of near the X point , with , in sharp contrast to systems with – band inversion such as MnBi2Te4 [49], where . Combined with the strong SOC from the heavy Ta atoms, this deep – inversion mechanism explains the large topological gaps summarized in Table 1.
@Ta | |||
@Ta | |||
@Fe |
Discussions—The realization of QAH state requires a delicate interplay between topology and magnetism. In FeTa, this balance is achieved through distinct yet cooperative mechanisms: the nontrivial band topology originates from SOC-driven inversion within Ta orbitals, while magnetism arises primarily from Fe electrons. Strong FM exchange between Fe atoms stabilizes local moments, inducing spin polarization on Ta sites and enabling polarized band inversion in the ferrimagnetic state. A crucial factor is the electron transfer from the less electronegative Ta to Fe, which establishes the desired Fe valence state. In contrast, LixFeSe is limited by chemically active and weakly bound Li+ ions, which cannot reach full intercalation () with current techniques [62], leading to disruptions in the topological electronic structure [41]. By comparison, Ta in FeTa forms strong and stable bonds with the ligands, thereby preserving the delicate electronic and magnetic environment essential for realizing the QAH state. This targeted substitution approach—simultaneously tuning topology, magnetism, and charge transfer—offers a general strategy for designing multifunctional materials that cohost topological, magnetic, and superconducting phases within a single compound.
FeTa constitutes the first example of a QAH insulator in a stoichiometric ferrimagnet. These materials simultaneously support spin transport and quantized edge charge currents, making them highly attractive for dissipationless electronics, spintronics, quantum computation, and magnetic tunnel junctions. The coexistence of topology and ferrimagnetism defines a promising new research direction. Owing to their inequivalent magnetic sublattices, ferrimagnets can potentially offer the combined advantages of both ferromagnets and antiferromagnets [79]: net magnetization is readily controllable via external fields, while spin dynamics remain fast, akin to AFM systems. It is intriguing to find compensated ferrimagnetic QAH insulators, in which the topological quantization efffect can be realized without any net magnetization and be controlled via fast magnetization reversal [80]. Moreover, the lattice constants of FeTa closely match those of Fe superconductors, facilitating the fabrication of FeTa/Fe heterostructures. The resulting superconducting proximity effect with reduced net magnetization offers a viable platform for realizing chiral topological superconductivity and Majorana fermions [10, 11, 12].
In summary, we identify a class of large-gap QAH insulators with Chern number , where magnetism and topology emerge cooperatively from orbital electrons alone. In contrast to earlier generations of magnetic TIs—such as magnetically doped TI and MnBi2Te4—the FeTa family uniquely combines structural simplicity, intrinsic magnetism, and topological nontriviality. We anticipate that these materials will serve as a next-generation platform for advancing research in topological quantum matter and spintronics.
Acknowledgements.
Acknowledgments—We thank J. Gao, T. Yu, J. Xiao, and Y. Zhang for valuable discussions. This work is supported by the Natural Science Foundation of China through Grants No. 12350404 and No. 12174066, the Innovation Program for Quantum Science and Technology through Grant No. 2021ZD0302600, the Science and Technology Commission of Shanghai Municipality under Grants No. 23JC1400600, No. 24LZ1400100 and No. 2019SHZDZX01. Y.J. acknowledges additional support from the China Postdoctoral Science Foundation under Grants No. GZC20240302 and No. 2024M760488.References
- Tokura et al. [2019] Y. Tokura, K. Yasuda, and A. Tsukazaki, Magnetic topological insulators, Nature Rev. Phys. 1, 126 (2019).
- Chang et al. [2023] C.-Z. Chang, C.-X. Liu, and A. H. MacDonald, Colloquium: Quantum anomalous hall effect, Rev. Mod. Phys. 95, 011002 (2023).
- Wang and Zhang [2017] J. Wang and S.-C. Zhang, Topological states of condensed matter, Nature Mat. 16, 1062 (2017).
- Bernevig et al. [2022] B. A. Bernevig, C. Felser, and H. Beidenkopf, Progress and prospects in magnetic topological materials, Nature 603, 41 (2022).
- Thouless et al. [1982] D. J. Thouless, M. Kohmoto, M. P. Nightingale, and M. den Nijs, Quantized hall conductance in a two-dimensional periodic potential, Phys. Rev. Lett. 49, 405 (1982).
- Haldane [1988] F. D. M. Haldane, Model for a quantum hall effect without landau levels: Condensed-matter realization of the ”parity anomaly”, Phys. Rev. Lett. 61, 2015 (1988).
- Halperin [1982] B. I. Halperin, Quantized hall conductance, current-carrying edge states, and the existence of extended states in a two-dimensional disordered potential, Phys. Rev. B 25, 2185 (1982).
- Fu and Kane [2009] L. Fu and C. L. Kane, Probing neutral majorana fermion edge modes with charge transport, Phys. Rev. Lett. 102, 216403 (2009).
- Akhmerov et al. [2009] A. R. Akhmerov, J. Nilsson, and C. W. J. Beenakker, Electrically detected interferometry of majorana fermions in a topological insulator, Phys. Rev. Lett. 102, 216404 (2009).
- Qi et al. [2010] X.-L. Qi, T. L. Hughes, and S.-C. Zhang, Chiral topological superconductor from the quantum hall state, Phys. Rev. B 82, 184516 (2010).
- Wang et al. [2015a] J. Wang, Q. Zhou, B. Lian, and S.-C. Zhang, Chiral topological superconductor and half-integer conductance plateau from quantum anomalous hall plateau transition, Phys. Rev. B 92, 064520 (2015a).
- Lian et al. [2018] B. Lian, X.-Q. Sun, A. Vaezi, X.-L. Qi, and S.-C. Zhang, Topological quantum computation based on chiral majorana fermions, Proc. Natl. Acad. Sci. USA 115, 10938 (2018).
- Qi et al. [2008] X.-L. Qi, T. L. Hughes, and S.-C. Zhang, Topological field theory of time-reversal invariant insulators, Phys. Rev. B 78, 195424 (2008).
- Essin et al. [2009] A. M. Essin, J. E. Moore, and D. Vanderbilt, Magnetoelectric polarizability and axion electrodynamics in crystalline insulators, Phys. Rev. Lett. 102, 146805 (2009).
- Morimoto et al. [2015] T. Morimoto, A. Furusaki, and N. Nagaosa, Topological magnetoelectric effects in thin films of topological insulators, Phys. Rev. B 92, 085113 (2015).
- Wang et al. [2015b] J. Wang, B. Lian, X.-L. Qi, and S.-C. Zhang, Quantized topological magnetoelectric effect of the zero-plateau quantum anomalous Hall state, Phys. Rev. B 92, 081107 (2015b).
- Mogi et al. [2017] M. Mogi, M. Kawamura, R. Yoshimi, A. Tsukazaki, Y. Kozuka, N. Shirakawa, K. S. Takahashi, M. Kawasaki, and Y. Tokura, A magnetic heterostructure of topological insulators as a candidate for an axion insulator, Nature Mater. 16, 516 (2017).
- Zhang and Zhang [2012] X. Zhang and S.-C. Zhang, Chiral interconnects based on topological insulators, Proc. SPIE Micro- Nanotechnol. Sens. Syst. Appl. IV 8373, 837309 (2012).
- Wang et al. [2013] J. Wang, B. Lian, H. Zhang, Y. Xu, and S.-C. Zhang, Quantum anomalous hall effect with higher plateaus, Phys. Rev. Lett. 111, 136801 (2013).
- Chang et al. [2013] C.-Z. Chang, J. Zhang, X. Feng, J. Shen, Z. Zhang, M. Guo, K. Li, Y. Ou, P. Wei, L.-L. Wang, Z.-Q. Ji, Y. Feng, S. Ji, X. Chen, J. Jia, X. Dai, Z. Fang, S.-C. Zhang, K. He, Y. Wang, L. Lu, X.-C. Ma, and Q.-K. Xue, Experimental Observation of the Quantum Anomalous Hall Effect in a Magnetic Topological Insulator, Science 340, 167 (2013).
- Chang et al. [2015] C.-Z. Chang, W. Zhao, D. Y. Kim, H. Zhang, B. A. Assaf, D. Heiman, S.-C. Zhang, C. Liu, M. H. W. Chan, and J. S. Moodera, High-precision realization of robust quantum anomalous hall state in a hard ferromagnetic topological insulator, Nature Mater. 14, 473 (2015).
- Mogi et al. [2015] M. Mogi, R. Yoshimi, A. Tsukazaki, K. Yasuda, Y. Kozuka, K. S. Takahashi, M. Kawasaki, and Y. Tokura, Magnetic modulation doping in topological insulators toward higher-temperature quantum anomalous hall effect, Appl. Phys. Lett. 107, 182401 (2015).
- Bestwick et al. [2015] A. J. Bestwick, E. J. Fox, X. Kou, L. Pan, K. L. Wang, and D. Goldhaber-Gordon, Precise quantization of the anomalous hall effect near zero magnetic field, Phys. Rev. Lett. 114, 187201 (2015).
- Watanabe et al. [2019] R. Watanabe, R. Yoshimi, M. Kawamura, M. Mogi, A. Tsukazaki, X. Z. Yu, K. Nakajima, K. S. Takahashi, M. Kawasaki, and Y. Tokura, Quantum anomalous hall effect driven by magnetic proximity coupling in all-telluride based heterostructure, Appl. Phys. Lett. 115, 102403 (2019).
- Deng et al. [2020] Y. Deng, Y. Yu, M. Z. Shi, Z. Guo, Z. Xu, J. Wang, X. H. Chen, and Y. Zhang, Quantum anomalous hall effect in intrinsic magnetic topological insulator mnbi2te4, Science 367, 895 (2020).
- Li et al. [2024] Y. Li, Y. Bai, Y. Feng, J. Luan, Z. Gao, Y. Chen, Y. Tong, R. Liu, S. K. Chong, K. L. Wang, et al., Reentrant quantum anomalous hall effect in molecular beam epitaxy-grown mnbi2te4 thin films (2024), arXiv:2401.11450 [cond-mat.mes-hall] .
- Lian et al. [2025] Z. Lian, Y. Wang, Y. Wang, W.-H. Dong, Y. Feng, Z. Dong, M. Ma, S. Yang, L. Xu, Y. Li, et al., Antiferromagnetic quantum anomalous hall effect under spin flips and flops, Nature 641, 70 (2025).
- Serlin et al. [2020] M. Serlin, C. L. Tschirhart, H. Polshyn, Y. Zhang, J. Zhu, K. Watanabe, T. Taniguchi, L. Balents, and A. F. Young, Intrinsic quantized anomalous hall effect in a moiré heterostructure, Science 367, 900 (2020).
- Li et al. [2021] T. Li, S. Jiang, B. Shen, Y. Zhang, L. Li, Z. Tao, T. Devakul, K. Watanabe, T. Taniguchi, L. Fu, J. Shan, and K. F. Mak, Quantum anomalous hall effect from intertwined moiré bands, Nature 600, 641 (2021).
- Park et al. [2023] H. Park, J. Cai, E. Anderson, Y. Zhang, J. Zhu, X. Liu, C. Wang, W. Holtzmann, C. Hu, Z. Liu, T. Taniguchi, K. Watanabe, J.-H. Chu, T. Cao, L. Fu, W. Yao, C.-Z. Chang, D. Cobden, D. Xiao, and X. Xu, Observation of fractionally quantized anomalous hall effect, Nature 622, 74 (2023).
- Xu et al. [2023] F. Xu, Z. Sun, T. Jia, C. Liu, C. Xu, C. Li, Y. Gu, K. Watanabe, T. Taniguchi, B. Tong, J. Jia, Z. Shi, S. Jiang, Y. Zhang, X. Liu, and T. Li, Observation of integer and fractional quantum anomalous hall effects in twisted bilayer , Phys. Rev. X 13, 031037 (2023).
- Lu et al. [2024] Z. Lu, T. Han, Y. Yao, A. P. Reddy, J. Yang, J. Seo, K. Watanabe, T. Taniguchi, L. Fu, and L. Ju, Fractional quantum anomalous hall effect in multilayer graphene, Nature 626, 759 (2024).
- Han et al. [2024] T. Han, Z. Lu, Y. Yao, J. Yang, J. Seo, C. Yoon, K. Watanabe, T. Taniguchi, L. Fu, F. Zhang, and L. Ju, Large quantum anomalous hall effect in spin-orbit proximitized rhombohedral graphene, Science 384, 647 (2024).
- Sha et al. [2024] Y. Sha, J. Zheng, K. Liu, H. Du, K. Watanabe, T. Taniguchi, J. Jia, Z. Shi, R. Zhong, and G. Chen, Observation of a chern insulator in crystalline abca-tetralayer graphene with spin-orbit coupling, Science 384, 414 (2024).
- Okazaki et al. [2022] Y. Okazaki, T. Oe, M. Kawamura, R. Yoshimi, S. Nakamura, S. Takada, M. Mogi, K. S. Takahashi, A. Tsukazaki, M. Kawasaki, Y. Tokura, and N.-H. Kaneko, Quantum anomalous hall effect with a permanent magnet defines a quantum resistance standard, Nature Phys. 18, 25 (2022).
- Patel et al. [2024] D. Patel, K. Fijalkowski, M. Kruskopf, N. Liu, M. Götz, E. Pesel, M. Jaime, M. Klement, S. Schreyeck, K. Brunner, et al., A zero external magnetic field quantum standard of resistance at the 10-9 level, Nature Electron. 7, 1111 (2024).
- Huáng et al. [2025] N. J. Huáng, J. L. Boland, K. M. Fijalkowski, C. Gould, T. Hesjedal, O. Kazakova, S. Kumar, and H. Scherer, Quantum anomalous hall effect for metrology, Appl. Phys. Lett. 126, 040501 (2025).
- You et al. [2019] J.-Y. You, Z. Zhang, B. Gu, and G. Su, Two-dimensional room-temperature ferromagnetic semiconductors with quantum anomalous hall effect, Phys. Rev. Applied 12, 024063 (2019).
- Sun et al. [2020a] J. Sun, X. Zhong, W. Cui, J. Shi, J. Hao, M. Xu, and Y. Li, The intrinsic magnetism, quantum anomalous hall effect and curie temperature in 2d transition metal trihalides, Phys. Chem. Chem. Phys. 22, 2429 (2020a).
- Sun et al. [2020b] Q. Sun, Y. Ma, and N. Kioussis, Two-dimensional dirac spin-gapless semiconductors with tunable perpendicular magnetic anisotropy and a robust quantum anomalous hall effect, Mater. Horiz. 7, 2071 (2020b).
- Li et al. [2020] Y. Li, J. Li, Y. Li, M. Ye, F. Zheng, Z. Zhang, J. Fu, W. Duan, and Y. Xu, High-temperature quantum anomalous hall insulators in lithium-decorated iron-based superconductor materials, Phys. Rev. Lett. 125, 086401 (2020).
- Li et al. [2022] Z. Li, Y. Han, and Z. Qiao, Chern number tunable quantum anomalous hall effect in monolayer transitional metal oxides via manipulating magnetization orientation, Phys. Rev. Lett. 129, 036801 (2022).
- Xuan et al. [2022] X. Xuan, Z. Zhang, C. Chen, and W. Guo, Robust quantum anomalous hall states in monolayer and few-layer tite, Nano Lett. 22, 5379 (2022).
- Jiang et al. [2023] Y. Jiang, H. Wang, and J. Wang, Large-gap quantum anomalous hall insulators in the (, rb, sr; =sb, bi, sn) class of compounds, Phys. Rev. B 108, 165122 (2023).
- Yao et al. [2024] Q. Yao, Y. Xue, B. Zhao, Y. Zhu, Z. Li, and Z. Yang, Orbital-selectivity-induced robust quantum anomalous hall effect in hund’s metals mgfep, Nano Lett. 24, 1563 (2024).
- Jiang et al. [2024] Y. Jiang, H. Wang, K. Bao, Z. Liu, and J. Wang, Monolayer v: A new family of quantum anomalous hall insulators, Phys. Rev. Lett. 132, 106602 (2024).
- Liu et al. [2008] C.-X. Liu, X.-L. Qi, X. Dai, Z. Fang, and S.-C. Zhang, Quantum anomalous hall effect in quantum wells, Phys. Rev. Lett. 101, 146802 (2008).
- Yu et al. [2010] R. Yu, W. Zhang, H.-J. Zhang, S.-C. Zhang, X. Dai, and Z. Fang, Quantized Anomalous Hall Effect in Magnetic Topological Insulators, Science 329, 61 (2010).
- Zhang et al. [2019] D. Zhang, M. Shi, T. Zhu, D. Xing, H. Zhang, and J. Wang, Topological axion states in the magnetic insulator with the quantized magnetoelectric effect, Phys. Rev. Lett. 122, 206401 (2019).
- Li et al. [2019] J. Li, Y. Li, S. Du, Z. Wang, B.-L. Gu, S.-C. Zhang, K. He, W. Duan, and Y. Xu, Intrinsic magnetic topological insulators in van der waals layered mnbi2te4-family materials, Sci. Adv. 5, eaaw5685 (2019).
- Otrokov et al. [2019] M. M. Otrokov, I. P. Rusinov, M. Blanco-Rey, M. Hoffmann, A. Y. Vyazovskaya, S. V. Eremeev, A. Ernst, P. M. Echenique, A. Arnau, and E. V. Chulkov, Unique thickness-dependent properties of the van der waals interlayer antiferromagnet films, Phys. Rev. Lett. 122, 107202 (2019).
- Chong et al. [2020] Y. X. Chong, X. Liu, R. Sharma, A. Kostin, G. Gu, K. Fujita, J. C. S. Davis, and P. O. Sprau, Severe dirac mass gap suppression in sb2te3-based quantum anomalous hall materials, Nano Lett. 20, 8001 (2020).
- Garnica et al. [2022] M. Garnica, M. M. Otrokov, P. C. Aguilar, I. I. Klimovskikh, D. Estyunin, Z. S. Aliev, I. R. Amiraslanov, N. A. Abdullayev, V. N. Zverev, M. B. Babanly, N. T. Mamedov, A. M. Shikin, A. Arnau, A. L. V. de Parga, E. V. Chulkov, and R. Miranda, Native point defects and their implications for the dirac point gap at mnbi2te4(0001), npj Quantum Mater. 7, 7 (2022).
- Liu and Wang [2025] Z. Liu and J. Wang, Layer-dependent quantum anomalous hall effect in rhombohedral graphene, Phys. Rev. B 111, L081111 (2025).
- Krukau et al. [2006] A. V. Krukau, O. A. Vydrov, A. F. Izmaylov, and G. E. Scuseria, Influence of the exchange screening parameter on the performance of screened hybrid functionals, J. Chem. Phys. 125, 224106 (2006).
- Kresse and Furthmüller [1996] G. Kresse and J. Furthmüller, Efficient iterative schemes for ab initio total-energy calculations using a plane-wave basis set, Phys. Rev. B 54, 11169 (1996).
- Dudarev et al. [1998] S. L. Dudarev, G. A. Botton, S. Y. Savrasov, C. J. Humphreys, and A. P. Sutton, Electron-energy-loss spectra and the structural stability of nickel oxide: An lsda+u study, Phys. Rev. B 57, 1505 (1998).
- Perdew et al. [1996] J. P. Perdew, K. Burke, and M. Ernzerhof, Generalized gradient approximation made simple, Phys. Rev. Lett. 77, 3865 (1996).
- Lei et al. [2017] B. Lei, N. Z. Wang, C. Shang, F. B. Meng, L. K. Ma, X. G. Luo, T. Wu, Z. Sun, Y. Wang, Z. Jiang, B. H. Mao, Z. Liu, Y. J. Yu, Y. B. Zhang, and X. H. Chen, Tuning phase transitions in fese thin flakes by field-effect transistor with solid ion conductor as the gate dielectric, Phys. Rev. B 95, 020503 (2017).
- Ying et al. [2018] T. P. Ying, M. X. Wang, X. X. Wu, Z. Y. Zhao, Z. Z. Zhang, B. Q. Song, Y. C. Li, B. Lei, Q. Li, Y. Yu, E. J. Cheng, Z. H. An, Y. Zhang, X. Y. Jia, W. Yang, X. H. Chen, and S. Y. Li, Discrete superconducting phases in fese-derived superconductors, Phys. Rev. Lett. 121, 207003 (2018).
- Kim et al. [2023] Y. H. Kim, S. Park, C. I. Kwon, S. Y. Kim, K. Watanabe, T. Taniguchi, and J. S. Kim, Two-dimensional multiband superconductivity of the optimally and uniformly li-intercalated fese nanoflakes, Curr. Appl. Phys. 46, 27 (2023).
- Hu et al. [2024] Y. Hu, F. Meng, H. Lei, Q.-K. Xue, and D. Zhang, Possible spin-polarized cooper pairing in high temperature fese superconductor (2024), arXiv:2405.10482 [cond-mat.supr-con] .
- Xu et al. [2016] G. Xu, B. Lian, P. Tang, X.-L. Qi, and S.-C. Zhang, Topological superconductivity on the surface of fe-based superconductors, Phys. Rev. Lett. 117, 047001 (2016).
- Wang et al. [2018] D. Wang, L. Kong, P. Fan, H. Chen, S. Zhu, W. Liu, L. Cao, Y. Sun, S. Du, J. Schneeloch, R. Zhong, G. Gu, L. Fu, H. Ding, and H.-J. Gao, Evidence for majorana bound states in an iron-based superconductor, Science 362, 333 (2018).
- Zhang et al. [2018] P. Zhang, K. Yaji, T. Hashimoto, Y. Ota, T. Kondo, K. Okazaki, Z. Wang, J. Wen, G. D. Gu, H. Ding, and S. Shin, Observation of topological superconductivity on the surface of an iron-based superconductor, Science 360, 182 (2018).
- Aroyo et al. [2006] M. I. Aroyo, J. M. Perezmato, C. Capillas, E. Kroumova, S. Ivantchev, G. Madariaga, A. Kirov, and H. Wondratschek, Bilbao crystallographic server: I. databases and crystallographic computing programs, Z. Krist. 221, 15 (2006).
- Kirov et al. [2006] A. Kirov, C. Capillas, J. Perez-Mato, and H. Wondratschek, Bilbao crystallographic server. ii. representations of crystallographic point groups and space groups, Acta Cryst. 62, 115 (2006).
- Perez-Mato et al. [2011] J. Perez-Mato, D. Orobengoa, E. Tasci, G. De la Flor Martin, and A. Kirov, Crystallography online: Bilbao crystallographic server, Bulg. Chem. Commun. 43, 183 (2011).
- Kruthoff et al. [2017] J. Kruthoff, J. de Boer, J. van Wezel, C. L. Kane, and R.-J. Slager, Topological classification of crystalline insulators through band structure combinatorics, Phys. Rev. X 7, 041069 (2017).
- Vergniory et al. [2017] M. G. Vergniory, L. Elcoro, Z. Wang, J. Cano, C. Felser, M. I. Aroyo, B. A. Bernevig, and B. Bradlyn, Graph theory data for topological quantum chemistry, Phys. Rev. E 96, 023310 (2017).
- Elcoro et al. [2017] L. Elcoro, B. Bradlyn, Z. Wang, M. G. Vergniory, J. Cano, C. Felser, B. Bernevig, D. Orobengoa, G. D. L. Flor, and M. Aroyo, Double crystallographic groups and their representations on the bilbao crystallographic server, J. Appl. Crystallogr 50, 1457 (2017).
- Bradlyn et al. [2017] B. Bradlyn, L. Elcoro, J. Cano, M. Vergniory, Z. Wang, C. Felser, M. Aroyo, and B. A. Bernevig, Topological quantum chemistry, Nature 547, 298 (2017).
- [73] See Supplemental Material for methods and technical details.
- Enayat et al. [2014] M. Enayat, Z. Sun, U. R. Singh, R. Aluru, S. Schmaus, A. Yaresko, Y. Liu, C. Lin, V. Tsurkan, A. Loidl, et al., Real-space imaging of the atomic-scale magnetic structure of Fe1+yTe, Science 345, 653 (2014).
- Goodenough [1963] J. B. Goodenough, Magnetism and the Chemical Bond (John Wiley And Sons, 1963).
- Spaldin [2012] N. A. Spaldin, Magnetic Materials: Fundamentals and Applications, 2nd Edition (Cambridge University Press, 2012).
- Khomskii [2004] D. I. Khomskii, Transition Metal Compounds (Cambridge University Press, 2004).
- Deng et al. [2018] Y. Deng, Y. Yu, Y. Song, J. Zhang, N. Z. Wang, Z. Sun, Y. Yi, Y. Z. Wu, S. Wu, J. Zhu, et al., Gate-tunable room-temperature ferromagnetism in two-dimensional fe3gete2, Nature 563, 94 (2018).
- Kim et al. [2022] S. K. Kim, G. S. Beach, K.-J. Lee, T. Ono, T. Rasing, and H. Yang, Ferrimagnetic spintronics, Nature Mater. 21, 24 (2022).
- Finley and Liu [2020] J. Finley and L. Liu, Spintronics with compensated ferrimagnets, Appl. Phys. Lett. 116, 110501 (2020).